Dorchester Company, on March 1, 2019 has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. On March 1, Dorchester started into production 16,500 units. At the end of the month there were 13,000 units completed and transferred into the Finished Goods Inventory. The ending WIP was 60% complete with respect to conversion. For the month of March the following costs were incurred and recorded in the WIP:

            Direct Material                      $31,000

            Direct Labor 32,000

            Factory Overhead 10,000

Dorchester uses the weighted-average process costing method. Use this information to determine the cost per equivalent unit of conversion for the month of March: (Round & enter final answers to the nearest cent.)
Answer #1

Answer

Physical unit

Particulars units
Beginning work in process 0
Started units 16,500
Total units 16,500
Units completed and transferred out 13,000
Ending work in process 3,500
Total units 16,500

Equivalent unit of conversion cost

= 13,000 + (3,500 × 60%) = 15,100 units

Cost per equivalent unit of material

= (32,000 + 10,000) / 15,100 = $2.78

$2.78 is the cost per equivalent unit of conversion cost under weighted average method.

The cost of conversion cost includes direct labour and factory overhead.

If YOU are like the answer please upvote.